OpenStudy (anonymous):

The perimeter of a rectangle is 132m the length is 6m more than twice the width find the dimensions. What is the width?

OpenStudy (anonymous):

W=20

OpenStudy (anonymous):

width =20 lenght = 46

OpenStudy (anonymous):

I did answer

OpenStudy (anonymous):

P=(l+W)*2 l=6+2W 132=(6+2W)+W)*2 132/2 =6+3W 66 -6 =3W 60/3 = W 20 = W L = 6+(2*20) L = 46

OpenStudy (anonymous):

check: P= (20 +46)*2 = 66*2 P = 132

OpenStudy (anonymous):

Thank you! I gave you a medal...here is another.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw